# From Context to EDUs: Faithful and Structured Context Compression via Elementary Discourse Unit Decomposition

### v3

@@ 3.1.1 Experiment Settings
We employ two complementary metrics to evaluate structural fidelity. Tree Edit Distance (TED) (ted_zhang1989simple) acts as a micro-level metric to measure structural dissimilarity by computing the minimum number of edit operations (insertion, deletion, substitution) required to transform the predicted tree into the ground truth, where a lower TED indicates more precise structural alignment. Complementarily, Document Level Accuracy (DLA) serves as a macro-level metric defined as $\mathrm{DLA}=\frac{\lvert D_{\text{match}}\rvert}{\lvert D_{\text{all}}\rvert}$, in which $D_{\text{match}}$ represents the count of documents where the decomposed structural backbone perfectly matches the ground truth. This constitutes a rigorous metric that requires zero structural errors.
@@
We compare our LingoEDU (built on Qwen3-4B (qwen3technicalreport)) against two categories of strong baselines. For all LLMs, we designed specific prompts to instruct them to output hierarchical JSON/Markdown structures: (1)Frontier LLMs: We evaluated SOTA models including GPT-4o/4.1 (gpt-4o), OpenAI o3/o4-mini (o3), Claude 3.7 Sonnet/4 (Claude4), DeepSeek-V3/R1 (deepseekai2025deepseekv3technicalreport), and Qwen3-235B (qwen3). All LLM results are averaged over three distinct runs. (2)Commercial Parsing APIs: We selected Jina Reader and Firecrawl, which are widely used for web-to-markdown conversion. We deployed test documents on a static server to allow URL-based access.
@@ 3.1.3 Ablation Studies
Table 2 validates the effectiveness of our design choices. First, the significant performance gap between “Indices Only” and our method (33.06% vs. 49.60% DLA) highlights that explicit text generation acts as a crucial semantic anchor for structural prediction. Second, the model exhibits remarkable data efficiency; even when scaled down to just 20% of the training data, it achieves a TED of 4.87 and retains over 91% of the full model’s accuracy.
@@
Finally, we investigate the impact of model scale on structural parsing capability, as shown in Table 3. Scaling the backbone from 1.7B to 4B yields clear improvements, reducing TED from 4.99 to 4.77. However, further scaling to 8B results in performance saturation: both TED and DLA slightly regress compared to the 4B model. This suggests that the 4B parameter range strikes an optimal balance for this specific task, whereas larger models may suffer from overfitting to the rigid output format without proportionally larger datasets.
@@
Table 4: Transposed results on LongBench. Datasets are grouped by task type (columns). $\Delta$ denotes the relative improvement of Ours over the Standard baseline for each specific backbone. Bold indicates best performance per backbone group; underlined indicates second-best.
@@ 3.2 Main Results on Downstream Long-Context Tasks
In this section, we address the second research question: Does structure-aware compression tangibly enhance performance for downstream tasks? We evaluate the utility of our EDU-based Context Compressor across two distinct scenarios: standard long-context benchmarks and complex open-domain Deep Search pipelines.
@@ 3.2.1 General Long-Context Understanding
We evaluate on LongBench (bai-etal-2024-longbench), a benchmark covering multi-document QA, summarization, and few-shot learning.
Baselines.
We compare three configurations: (1) Standard, which feeds the full (or truncated) original context directly to the LLM; (2) Self-Sum, which utilizes the LLM itself to generate abstractive summaries of the context prior to processing; and (3) Glyph (cheng2025glyph), which applies an implicit compression baseline to serve as a reference for latent-based methods.
@@
Table 5: Ablation Study on Ranking Models. All methods use GPT-4.1 as the final generator. We compare No Selection, Random, BM25, LLM Self-Selection, and Our Dedicated Reranker. Bold denotes the best performance.
@@
Table 6: Ablation study of the LingoEDU module on Deep Search benchmarks. Accuracy scores (%) are reported. Base: Standard Deep Search without compression; Self-Sum: Query-focused summarization; Ours (LingoEDU): Structural decomposition. $\Delta$ denotes the relative improvement of Ours over the Base baseline.
@@ 3.2.2 Impact on Deep Search
Critically, our method remains additive even for the most advanced models like GPT-5 and Gemini 3 Pro (w/o Deep Think)(Table 6, rows 4-6). This indicates that even as model capacity grows, handling unstructured noise and long-context reasoning remain bottlenecks. Our explicit compression provides a complementary signal that enhances the internal processing of SOTA LLMs.
@@ Appendix C Experiment Details
C.1 Detailed Model Specifications
@@
C.2 Robustness Analysis Across Context Lengths
@@ 3. Answering Phase
C.4 Impact of Retrieval Granularity on Compression and Performance

### v4

@@ 3.1.1 Experiment Settings
We employ two complementary metrics to evaluate structural fidelity. Tree Edit Distance (TED) (ted_zhang1989simple) acts as a micro-level metric to measure structural dissimilarity by computing the minimum number of edit operations (insertion, deletion, substitution) required to transform the predicted tree into the ground truth, where a lower TED indicates more precise structural alignment. Complementarily, Document Level Accuracy (DLA) serves as a macro-level metric defined as $\mathrm{DLA}=\frac{\lvert D_{\text{match}}\rvert}{\lvert D_{\text{all}}\rvert}$, in which $D_{\text{match}}$ represents the count of documents where the decomposed structural backbone perfectly matches the ground truth. This rigorous metric requires zero structural errors.
@@
We compare our LingoEDU (built on Qwen3-4B (qwen3technicalreport)) against two categories of strong baselines. For all LLMs, we designed specific prompts to instruct them to output hierarchical JSON/Markdown structures: (1)Frontier LLMs: We evaluated SOTA models including GPT-4o/4.1 (gpt-4o), OpenAI o3/o4-mini (o3), Claude 3.7 Sonnet/4 Sonnet (Claude4), DeepSeek-V3/R1 (deepseekai2025deepseekv3technicalreport), and Qwen3-235B (qwen3). All LLM results are averaged over three distinct runs. (2)Commercial Parsing APIs: We selected Jina Reader and Firecrawl, which are widely used for web-to-markdown conversion. We deployed test documents on a static server to allow URL-based access.
@@ 3.1.3 Ablation Studies
Table 2 validates the effectiveness of our design choices. First, the significant performance gap between “Indices Only” and our method highlights that explicit text generation acts as a crucial semantic anchor for structural prediction. Second, the model exhibits remarkable data efficiency; even when scaled down to just 20% of the training data, it achieves a TED of 4.87 and retains over 91% of the full model’s accuracy.
@@
Finally, we investigate the impact of model scale on structural parsing capability, as shown in Table 3. Scaling the backbone from 1.7B to 4B yields clear improvements, reducing TED from 4.99 to 4.77. However, further scaling to 8B results in performance saturation: both TED and DLA regress compared to the 4B model. This suggests that the 4B parameter range strikes an optimal balance for this task, whereas larger models may suffer from overfitting to the rigid output format without proportionally larger datasets.
@@
Table 4: Results on LongBench. Datasets are grouped by task type (columns). $\Delta$ denotes the relative improvement of Ours over the Standard baseline for each specific backbone. Bold indicates best performance per backbone group; underlined indicates second-best.
@@ 3.2 Main Results on Downstream Long-Context Tasks
In this section, we address the second research question: Does structure-aware compression tangibly enhance performance for downstream tasks? We evaluate the utility of our EDU-based Context Compressor across two distinct scenarios: standard long-context benchmarks and complex open-domain Deep Search.
@@ 3.2.1 General Long-Context Understanding
We evaluate our LingEDU on LongBench (bai-etal-2024-longbench), a benchmark covering multi-document QA, summarization, and few-shot learning.
Baselines. We compare three configurations: (1) Standard, which feeds the full original context directly to the LLM; (2) Self-Sum, which utilizes the LLM itself to generate abstractive summaries of the context prior to processing; (3) Glyph (cheng2025glyph), which applies an implicit compression baseline to serve as a reference for latent-based methods; and (4) C3 (liu2025context), which employs a small LLM to aggressively compress long contexts into compact latent representations before decoding with a LLM.
@@
Table 5: Ablation study on ranking models. All methods use GPT-4.1 as the final generator. We compare No Selection, Random, BM25, LLM Self-Selection, and Our Dedicated Reranker.
@@
Table 6: Ablation study of the LingoEDU module on Deep Search. Accuracy scores (%) are reported. Base: Standard Deep Search without compression; Self-Sum: Query-focused summarization; Ours (LingoEDU): Structural decomposition. $\Delta$ denotes the relative improvement of Ours over the Base baseline.
@@ 3.2.2 Impact on Deep Search
Critically, our method remains additive even for the most advanced models like GPT-5 and Gemini 3 Pro (w/o Deep Think). This indicates that even as model capacity grows, handling unstructured noise and long-context reasoning remain bottlenecks. Our explicit compression provides a complementary signal that enhances the internal processing of SOTA LLMs.
@@ Appendix C Experiment Details
C.1 Train Details
Hyperparameter Value
Model Configuration
Base Model Qwen3-4B
Max Sequence Length 32,768
RoPE Base 1,000,000
Precision bf16
Optimization
Optimizer AdamW
Optimizer Params $\beta_{1}=0.9,\beta_{2}=0.95$
Peak Learning Rate $1\times 10^{-5}$
Min Learning Rate $1\times 10^{-6}$
LR Scheduler Cosine
Warmup Ratio 0.1
Weight Decay 0.1
Gradient Clipping 0.5
Batching & Parallelism
Global Batch Size 128
Training Iterations 1,296
Tensor Parallelism (TP) 4
Sequence Parallelism True
Table 8: Hyperparameters and configuration used for training the LingEDU.
We present the detailed training configuration in Table 8. The model is fine-tuned based on the Qwen3-4B architecture. To accommodate the long-context requirement, we extend the maximum sequence length to 32,768 tokens and adjust the Rotary Positional Embedding (RoPE) base frequency to $1,000,000$. We utilize the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\text{e-}8$. A weight decay of $0.1$ and a gradient clipping threshold of $1.0$ are applied to stabilize training. The learning rate follows a cosine decay schedule, starting from a peak of $1\text{e-}5$ and decaying to a minimum of $1\text{e-}6$, with a linear warmup phase covering the first 10% of the training steps.
C.2 Detailed Model Specifications
@@
C.3 Robustness Analysis Across Context Lengths
@@ 3. Answering Phase
C.5 Impact of Retrieval Granularity on Compression and Performance